OpenStudy (anonymous):

A student attempts to solve for a in the equation below. Which statement best applies to the work shown below?

OpenStudy (anonymous):

-3a-1/2=1 -3a*(-1/3)-1/2=1*(-1/3) a-1/2+1/2=1/3+1/2 a=1/6

OpenStudy (anonymous):

A. The student incorrectly applied the multiplication property of equality before the addition property of equality. B. The student did not apply the addition property of equality to isolate a. C. The student did not apply the multiplication property of equality to isolate a. D. The student correctly solved the two-step equation.